A particle moving with a non zero velocityand constant acceleration on straightline travels 15m in first 3s and 33m in next 3s in the same direction find initial velocity and acceleration of particle.
 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Varsha
 

using the eq _-- s= ut +1/2 a t^2

15=3u+9/2a----5=u+3/2a ........1

next 33 m in 3s,

tot dist=15+33=48

tot tme=3+3=6

48=6u+18a--- 8=u+3a.......2

equating both we get,

3=3/2a == a = 2m/s^2

u=2m/s.....